### v1

# Database Views as Explanations for Relational Deep Learning

Abstract.
In recent years, there has been significant progress in the development of deep learning models over relational databases, including architectures based on heterogeneous graph neural networks (hetero-GNNs) and heterogeneous graph transformers. In effect, such architectures state how the database records and links (e.g., foreign-key references) translate into a large, complex numerical expression, involving numerous learnable parameters. This complexity makes it hard to explain, in human-understandable terms, how a model uses the available data to arrive at a given prediction.
We present a novel framework for explaining machine-learning models over relational databases, where explanations are view definitions that highlight focused parts of the database that mostly contribute to the model’s prediction. We establish such global abductive explanations by adapting the classic notion of determinacy by Nash, Segoufin, and Vianu (2010). In addition to tuning the tradeoff between determinacy and conciseness, the framework allows controlling the level of granularity by adopting different fragments of view definitions, such as ones highlighting whole columns, foreign keys between tables, relevant groups of tuples, and so on.
We investigate the realization of the framework in the case of hetero-GNNs. We develop heuristic algorithms that avoid the exhaustive search over the space of all databases. We propose techniques that are model-agnostic, and others that are tailored to hetero-GNNs via the notion of learnable masking. Our approach is evaluated through an extensive empirical study on the RelBench collection, covering a variety of domains and different record-level tasks. The results demonstrate the usefulness of the proposed explanations, as well as the efficiency of their generation.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at URL_TO_YOUR_ARTIFACTS.
1. Introduction
Machine learning over relations was traditionally done through a reduction to learning over unstructured data: flatten the data into a feature matrix, with a collection of predefined queries that produce features for every prediction instance, and then learn a flat model (Perlich and Provost, 2003; Kimelfeld and Ré, 2018; Bakshi et al., 2021). Progress in artificial neural networks, and specifically Graph Neural Networks (GNNs) (Wu et al., 2020) and graph transformers (Müller et al., 2024) changed this practice: the best performing models now deploy prediction models directly from the structured data without ad-hoc feature extraction(Chen et al., 2025; Wang et al., 2024).
Specifically, deep learning is commonly applied to relational databases by translating them into a graph representation: nodes represent database entities (usually tuples), and edges represent relationships (usually foreign-primary key constraints). GNNs or graph transformers are then trained on these graphs (Fey et al., 2024; Cappuzzo et al., 2020; Tönshoff et al., 2023; Schlichtkrull and Søgaard, 2017). To handle the varying types of nodes and edges determined by the database schema, heterogeneous graph models are used (Zhao et al., 2021; Wang et al., 2019; Chen et al., 2025; Wang et al., 2024). This creates a two-paradigm pipeline: the relational database defines the graph structure, while the (graph-based) machine learning library learns the model. At inference time, new data is exported to the graph representation, predictions are computed from the learned model, and results are reintegrated into the database.
Figure 1. An explanation view defines a subdatabase (or more generally, a derived view). The rest of the database gets randomly perturbed. As long as the view is kept the same, a trained black-box model gives similar predictions over the perturbed database as the original database. We say that the view soft determines the prediction.
The neural network model induced by the above graph-learning approach over the database can be viewed as a complex numerical function constructed from the database’s tuples and values, and parameterized by a large number of learned weights. Consequently, as in other modalities of deep learning, the model’s internal behavior is inherently opaque and cannot be readily understood through direct inspection. To address this, the field of Explainable Artificial Intelligence (XAI) provides tools that shed light on different aspects of the model (Angelov et al., 2021; Arenas, 2024; Marques-Silva, 2023; Darwiche, 2023). These include local explanations, which account for the model’s output on a specific instance, and global explanations, which aim to characterize the model’s behavior as a whole. Explanations may be contrastive (sometimes called counterfactual (Barceló et al., 2025)), identifying hypothetical changes that would alter the model’s outcome, or abductive, pointing to focused parts of the input that suffice to determine the outcome (Marques-Silva, 2023).
In this work, we study global abductive explanations for deep relational models. Our focus is on settings where the underlying schema consists of multiple interrelated tables with numerous attributes—a common scenario in real-world databases. We aim to identify concise, informative subsets of the data that best account for the model’s predictions. Specifically, we address the question: Which parts of the database does the model rely on most when making predictions? Hence, our goal is to provide global insights into the model’s behavior by uncovering the database fragments that are most influential across predictions.
Example 1.
To illustrate the nature of our explanations, consider the “rel-trial” database (Initiative et al., 2019) from the RelBench collection (Robinson et al., 2024). Suppose we wish to explain a binary classifier that solves the following task: “Predict whether a trial will achieve its primary outcome.” Here, a trial corresponds to a record in the relation Studies. Our approach identifies only $\sim 10\%$ out of $140$ columns together with the $10$ relevant primary-foreign key pairs, which capture the schema components that largely determine the model’s output. In particular, our explanation highlights the attributes name, city, and country from the Facilities relation, which is linked to Studies via the intermediate relation Facilities_studies. This indicates that the names and locations of the utilized facilities highly influence the success or failure of the study. Interestingly, no attributes from the Outcomes relation, e.g., the outcome type or description, are included in the explanation, indicating that they are not actually as useful for the model to predict whether this outcome is achieved. Our explanations further show that specific study designs (i.e., records in the Designs relation) that utilize simple interventions are important for the model’s prediction. A detailed case study follows in Section 6.3. To verify that the schema components identified by our explanation capture the most important parts of the database, we train a smaller model using only those components, which achieves similar performance (ROC_AUC drops only slightly from $71.28$ to $69.78$). This is still high performance for the task (see (Robinson et al., 2024) for reference), while reducing the database by $\sim 80\%$.
To establish insights such as those in Example 1, we propose a general framework in which explanations are expressed as view definitions written in SQL (or any other query language of choice). This grounds explanations in a representation that is familiar to database users, in contrast to common alternatives that require machine learning expertise to interpret (Räuker et al., 2023). For graph-based models, this would require knowledge of the graph modeling details or even the inner workings of GNNs (Yuan et al., 2022).
To this aim, we deploy the classical notion of determinacy, introduced by Nash, Segoufin, and Vianu (Nash et al., 2010), that captures in a fundamental way when a collection of database views is enough to determine the result of a query. For reasons of usability, flexibility, and practical implementation, we propose a soft and statistical adaptation of this concept. Roughly speaking, the definition we propose identifies an explanation as accurate if the model is likely to have similar predictions on our database as it would have on alternative databases, as long as they agree on the views. Figure 1 shows a visual illustration of the concept.
In addition to accuracy, we seek concise explanations that are easy to digest. Clearly, the verbosity of an explanation depends on both the number of views and the complexity of each view. Our framework naturally supports the tradeoff between accuracy and conciseness, allowing users to balance interpretability with fidelity to the model’s behavior.
Nevertheless, the theoretical definition that we propose for explanations does not lend itself directly to an implementation. We develop heuristic techniques for the case of GNNs and important classes of views: projections, joins over foreign keys, and selections (filters). These heuristics are generally based on learning masked variations of the GNN at hand, where we place the masks in components of the GNN that correspond to the different kinds of views.
Lastly, we present an empirical evaluation of our techniques on the RelBench collection of databases and node-level tasks (Robinson et al., 2024). In particular, the experiments show the tradeoff between the accuracy and conciseness of the explanations. As a side benefit, they also show that, for several of the RelBench tasks, we can achieve models of similar quality to the ones publicly available, while using only a fraction of the database.
Our main contributions. 1 We introduce a model-agnostic framework for generating relational explanations of learned model behavior over relational databases, based on a statistical adaptation of determinacy. The framework uses relational queries, supports user-controlled granularity, and does not rely on internal model representations. 2 We instantiate the framework for relational deep learning using GNNs, and develop an efficient masking-based heuristic to generate explanations without explicitly estimating soft determinacy. 3 We conduct extensive experiments on the RelBench benchmark across several databases and tasks, proposing evaluation baselines and metrics. Our method consistently produces high-quality explanations at significantly lower computational cost.
2. Preliminaries
In this section, we define the concepts and notation that we use throughout the paper.
2.1. Relational Database Concepts
We begin with database concepts and terminology.
Databases
A relation schema $\mathcal{R}$ is associated with a sequence $\mathsf{att}(\mathcal{R})=(A_{1},\dots,A_{k})$ of distinct attributes, and a key signature $\mathsf{key}(\mathcal{R})=(B_{1},\dots,B_{\ell})$, so that $\mathsf{key}(\mathcal{R})$ is a subsequence of $\mathsf{att}(\mathcal{R})$. Each attribute $A_{i}$ is associated with a (finite or infinite) domain, denoted $\mathsf{dom}(A_{i})$. If $\mathord{\mathbf{C}}=(C_{1},\dots,C_{m})$ is a sequence of attributes of $R$, then we denote by $\mathsf{dom}(\mathord{\mathbf{C}})$ the set $\mathsf{dom}(C_{1})\times\dots\times\mathsf{dom}(C_{m})$. A tuple $t$ over a relation schema $\mathcal{R}$ is a member of $\mathsf{dom}(\mathsf{att}(R))$. For $i=1,\dots,k$, we denote by $t[A_{i}]$ the $i$-th value in $t$ (corresponding to the attribute $A_{i}$). If $\mathord{\mathbf{C}}=(C_{1},\dots,C_{m})$ is a sequence of attributes from $\mathsf{att}(\mathcal{R})$, then we denote by $t[\mathord{\mathbf{C}}]$ the tuple $(t[C_{1}],\dots,t[C_{m}])$.
A relation over the relation schema $\mathcal{R}$ is a finite set of tuples over $\mathcal{R}$ so that $t_{1}[\mathsf{key}(\mathcal{R})]\neq t_{2}[\mathsf{key}(\mathcal{R})]$ for every two distinct tuples $t_{1}\neq t_{2}$. If $r$ is a relation over $\mathcal{R}$ and $\mathord{\mathbf{C}}=(C_{1},\dots,C_{m})$ is a sequence of attributes from $\mathsf{att}(\mathcal{R})$, then $r[\mathord{\mathbf{C}}]$ denotes the relation $\mathord{\{t[\mathord{\mathbf{C}}]\mid t\in r\}}$. We denote with $|r|$ the number of tuples in $r$.
A database schema $\mathcal{S}$ is associated with a set $\mathsf{rel}(\mathcal{S})$ of relation names, and it maps every relation name $R\in\mathsf{rel}(\mathcal{S})$ to a relation schema $\mathcal{S}(R)$. To ease the notation, we may write $\mathsf{att}(R)$ and $\mathsf{key}(R)$ as a shorthand notation for $\mathsf{att}(\mathcal{S}(R))$ and $\mathsf{key}(\mathcal{S}(R))$, respectively. In addition, a database schema $\mathcal{S}$ includes a set $\mathsf{FK}(\mathcal{S})$ of foreign-key constraints, or FK for short, where an FK is an expression of the form $R[\mathord{\mathbf{C}}]\sqsubseteq\mathsf{key}(S)$, where $R$ and $S$ are relation names in $\mathsf{rel}(\mathcal{S}$), the sequence $\mathord{\mathbf{C}}$ consists of attributes from $\mathsf{att}(R)$, the sequences $\mathord{\mathbf{C}}$ and $\mathsf{key}(S)$ have the same length, and $\mathsf{dom}(\mathord{\mathbf{C}})\subseteq\mathsf{dom}(\mathsf{key}(S))$. A database $D$ over $\mathcal{S}$ maps every relation name $R\in\mathsf{rel}(\mathcal{S})$ to a relation $D(R)$ over $\mathcal{S}(R)$, so that every key constraint and every constraint in $\mathsf{FK}(\mathcal{S})$ is satisfied: the FK $R[\mathord{\mathbf{C}}]\sqsubseteq\mathsf{key}(S)$ is satisfied if $D(R)[\mathord{\mathbf{C}}]\subseteq D(S)[\mathsf{key}(S)]$.
Relational queries and predictive models
Let $\mathcal{S}$ be a database schema. A query $Q$ over $\mathcal{S}$ is associated with a relation schema, which we consistently denote by $\mathcal{R}_{Q}$, and it maps every database $D$ over $\mathcal{S}$ to a relation $Q(D)$ over $\mathcal{R}_{Q}$.
If $R\in\mathsf{rel}(\mathcal{S})$ is a relation name, then an $R$-model is a query (possibly very complicated and with learned parameters) over $\mathcal{S}$ whose purpose is to make a prediction (e.g., classification or regression) for every tuple in $R$. Formally, an $R$-model (of dimension $d$) is a query $M$ with the following properties:
  • existence of prediction attribute $A$: $\mathsf{att}(\mathcal{R}_{M})=(\mathsf{key}(R),A)$, where $A$ is an attribute with $\mathsf{dom}(A)=\mathbb{R}^{d}$, where $\mathbb{R}$ stand for the set of real numbers.
  • key signature agreement: $\mathsf{key}(\mathcal{R}_{M})=\mathsf{key}(R)$.
  • instance-prediction relation correspondence: For every database $D$, $M(D)[\mathsf{key}(R)]=D(R)[\mathsf{key}(R)]$. In words, if $r=M(D)$ is the relation resulting from applying $M$ to $D$, then $r$ has the exact same key values as the $R$-relation of $D$.
We refer to each tuple in $D(R)[\mathsf{key}(R)]$ (notice the projection on only the key attributes) as an instance (to be classified by $M$).
Let $M$ be an $R$-model and $D$ a database, both over the schema $\mathcal{S}$. Let $s$ be an instance (i.e., a key tuple of $D(R)$) to which we apply the model We denote by $M_{D}(s)$ the unique prediction of $M$ for $s$, that is, the value $t[A]\in\mathbb{R}^{d}$ of the unique tuple $t\in M(D)$ with $t[\mathsf{key}(R)]=s$.
Determinacy
We briefly recall the concept of determinacy by Nash et al. (2010). Let $\mathcal{S}$ be a schema, let $V_{1}$, …, $V_{m}$ be queries over $\mathcal{S}$, referred to as views, and let $Q$ be a query over $\mathcal{S}$. We say that $\mathord{\{V_{1},\dots,V_{m}\}}$ determines $Q$ if for all databases $D_{1}$ and $D_{2}$, if $V_{i}(D_{1})=V_{i}(D_{2})$ for every $i=1,\dots,m$, then $Q(D_{1})=Q(D_{2})$. Hence, knowing the result of every view suffices to determine the result of $Q$ without even looking at the database itself.
2.2. Heterogeneous Graphs
Next, we present the concepts and notation related to machine learning over heterogeneous graphs.
Hetero-graphs
A heterogeneous featured graph, or just hetero-graph for short, is a quadruple $\mathord{\mathcal{G}}=(V,E,\tau,\mathord{\mathbf{X}})$ where:
  • $V$ is a finite set of nodes;
  • $E$ is a set of typed directed edges, that is, triples $(u,\lambda,v)$ where $u$ and $v$ are nodes in $V$ and $\lambda$ is an edge type;
  • $\tau(\cdot)$ is a node-type function, mapping every node $v\in V$ to a type $\tau(v)$; and
  • $\mathord{\mathbf{X}}$ is a node-feature function that maps every node $v\in V$ to a vector $\mathord{\mathbf{X}}_{v}\in\mathbb{R}^{d_{\tau(v)}}$ for some predefined feature dimension $d_{\tau(v)}$ for each node type.
Every edge type $\lambda$ in $G$ is associated with a source type ${\mathsf{src}}(\lambda)$ and a target type ${\mathsf{tgt}}(\lambda)$, and we require all edges $e=(u,\lambda,v)$ to be consistent in that $\tau(u)={\mathsf{src}}(\lambda)$ and $\tau(v)={\mathsf{tgt}}(\lambda)$. We denote by $\tau(V)$ the set $\mathord{\{\tau(v)\mid v\in V\}}$. If $\alpha\in\tau(V)$, then we denote by $V[\alpha]$ the set of nodes $v\in V$ with $\tau(v)=\alpha$.
We denote by $\mathcal{N}^{\mathsf{in}}(v)$ (resp., $\mathcal{N}^{\mathsf{out}}(v)$) the set of incoming (resp., outgoing) neighbors of $v$, that is, the set $\mathord{\{u\mid\exists\lambda[(u,\lambda,v)\in E]\}}$ (resp., $\mathord{\{u\mid\exists\lambda[(v,\lambda,u)\in E]\}}$). If $\lambda$ is an edge type, then we denote by $\mathcal{N}^{\mathsf{in}}_{\lambda}(v)$ the set of nodes $u\in\mathcal{N}^{\mathsf{in}}(v)$ such that $\tau(u,\lambda,v)\in E$; analogously, we denote by $\mathcal{N}^{\mathsf{out}}_{\lambda}(v)$ the set of nodes $u\in\mathcal{N}^{\mathsf{out}}(v)$ with $\tau(v,\lambda,u)\in E$.
Node-centric models
Let $\mathord{\mathcal{G}}=(V,E,\tau,\mathord{\mathbf{X}})$ be a hetero-graph, and let $\alpha\in\tau(V)$ be a type, which we refer to as a prediction entity type. A $\alpha$-model (of dimension $d$) is a function $M:V[\alpha]\rightarrow\mathbb{R}^{d}$, mapping every $\alpha$-node $v$ to a predicted vector $M(v)$. Note that the dimension $d$ depends on the task (e.g., it is one in binary classification and regression, and $k$ in $k$-class classification), and is not necessarily the same as $d_{\alpha}$.
2.3. Heterogeneous Graph Neural Networks
A GNN defines a model over a graph by applying message passing to produce a vector representation of each node; in each step, called layer, the vector of a node is updated by aggregating the vectors of its neighbors, referred to as messages. A hetero-GNN operates over a hetero-graph similarly, except that messages of each edge type are aggregated separately. We give a more precise definition of a typical hetero-GNN next. We note that our definition is slightly more general and detailed than common definitions (e.g., (Barceló et al., 2022)), stressing the role of labels and directionality in the GNN.
Let $\mathord{\mathcal{G}}=(V,E,\tau,\mathord{\mathbf{X}})$ be a hetero-graph and $\alpha\in\tau(V)$. A hetero-GNN with $L$ layers defines an $\alpha$-model using the functions $\mathord{\mathbf{h}}^{(\ell)}$ and $\mathcal{A}^{(\ell)}$, for $\ell=0,\dots,L$, constructed inductively as follows. First, define $\mathord{\mathbf{h}}^{0}(v)=\mathord{\mathbf{X}}_{v}$ for all $v\in V$. For $\ell=0,\dots,L-1$, define
$\mathcal{A}^{(\ell)}(v,\lambda,o)=\mathsf{AGG}(\mathord{\{\mskip-5.0mu\{\mathord{\mathbf{h}}^{(\ell)}(u)\mid u\in\mathcal{N}^{o}_{\lambda}(v)\}\mskip-5.0mu\}})$
for every node $v$, edge label $\lambda$ and orientation $o\in\mathord{\{{\mathsf{in}},{\mathsf{out}}\}}$ such that $\mathcal{N}^{o}_{\lambda}(v)$ is defined.
Here, $\mathsf{AGG}$ is an aggregation function such as average or max, and $\mathord{\{\mskip-5.0mu\{\cdot\}\mskip-5.0mu\}}$ is the bag (multiset) notation. Next, for $\ell>0$ we define
(1) $\begin{split}\mathord{\mathbf{h}}^{(\ell)}(v)=\mathsf{UPD}(&\mathord{\mathbf{h}}^{(\ell-1)}(v),\\ &\mathcal{A}^{(\ell-1)}(v,\lambda_{1},o_{1}),\dots,\\ &\mathcal{A}^{(\ell-1)}(v,\lambda_{q},o_{q}))\end{split}$
for some ordering $(\lambda_{1},o_{1}),\dots,(\lambda_{q},o_{q})$ of all pairs $(\lambda,o)$ where $\mathcal{N}^{o}_{\lambda}(v)$ is defined. Here, $\mathsf{UPD}$ is an update function, which is realized as a neural model. A typical example of model architecture for hetero-GNNs (Schlichtkrull et al., 2018) involves a linear combination followed by an activation function (e.g., ReLU). The model $M$ is given by
$M(v)=\mathsf{MLP}(\mathord{\mathbf{h}}^{(L)}(v))$
where $\mathsf{MLP}$ is a multilayer-perception function that transforms the last layer into the actual prediction. The learnable parameters are incorporated in the functions $\mathsf{AGG}$, $\mathsf{UPD}$, and $\mathrm{MLP}$.
2.4. Relational Learning via Heterogeneous Graphs
A common way of applying deep learning over relational databases is to convert the database $D$ into a hetero-graph $\mathord{\mathcal{G}}_{D}$, and then learn and apply a hetero-GNN over $\mathord{\mathcal{G}}_{D}$ (Cvitkovic, 2019; Cappuzzo et al., 2024; Fey et al., 2024; Wang et al., 2024). Next, we describe this process more precisely for a typical conversion, known as the relational graph due to the r2n (i.e., row-to-node) method (Wang et al., 2024).
Let $\mathcal{S}$ be a schema, and let $D$ be a database over $\mathcal{S}$. The graph $\mathord{\mathcal{G}}_{D}=(V,E,\tau,\mathord{\mathbf{X}})$ is defined as follows. The set $V$ of nodes consists of the pairs $(R,t)$, where $R$ is a relation name and $t$ is tuple in $D(R)$:
$V=\mathord{\{(R,t)\mid R\in\mathsf{rel}(\mathcal{S})\land t\in D(R)\}}\,.$
For a node $v=(R,t)$, we define $\tau(v)=R$, that is, the type of a node is the name of its relation. There is a directed edge $e=(v,\varphi,u)$ from a node $v=(R,t)$ to a node $s=(S,t^{\prime})$ whenever the two are connected by a reference via an FK $\varphi$ of the form $R[\mathord{\mathbf{C}}]\sqsubseteq\mathsf{key}(S)$; in particular, the type of the edge is the FK that it corresponds to. Finally, the feature vector $\mathord{\mathbf{X}}$ is defined as follows. Let $R$ be a relation name with $\mathsf{att}(R)=(A_{1},\dots,A_{k})$. For each attribute $A$ we assume a trainable encoder $\mathsf{Enc}_{R.A}:\mathsf{dom}(A)\rightarrow\mathbb{R}^{d_{R.A}}$ for some predefined dimension $d_{R.A}$. In addition, for each relation name $R$ we assume a trainable encoder $\mathsf{Enc}_{R}:\mathbb{R}^{d_{R_{\mathit{enc}}}}\rightarrow\mathbb{R}^{d_{R}}$, for some dimension $d_{R}$ and $d_{R_{\mathit{enc}}}=\sum_{i=1}^{k}d_{R.A_{i}}$. For a node $v=(R,t)$ we then define:
(2) $\mathord{\mathbf{X}}_{v}=\mathsf{Enc}_{R}\left(\mathsf{Enc}_{R.A_{1}}(t[A_{1}])\oplus\cdots\oplus\mathrm{Enc}_{R.A_{k}}(t[A_{k}])\right)$
where $\oplus$ stands for vector concatenation. Notice that this formulation allows us to exclude some attributes $A$ of a relation $R$ from the encoding, simply by setting $\mathrm{Enc}_{R.A_{k}}$ to map every value to the empty vector. For example, in our implementation, we exclude key and foreign-key attributes from the features of a record.
Finally, let $R$ be a relation name in $\mathsf{rel}(\mathcal{S})$. An $R$-model $M^{\prime}$ over $\mathord{\mathcal{G}}_{D}$ defines an $R$-model $M$ over the database $D$ in a straightforward way: For a tuple $t\in D(R)$, we define
$M(t[\mathsf{key}(R)])=M^{\prime}(v)$
where $v$ is the node $(R,t)$ in the node set $V[R]$.
3. RDL Explanations via Soft Determinacy
In this section, we introduce our framework, where we leverage the concept of view determinacy to produce explanations.
3.1. Explanation Views
Let $\mathcal{S}$ be a schema with a relation name $R$, and let $M$ be an $R$-model. Let $D$ be a database over $\mathcal{S}$. We think of $M$ as a black-box complex model, involving a considerable number of computational components and learned parameters. By an explanation of $M$’s behavior on $D$, we refer to a collection of views $V$ (called explanation views), each indicating a focused component of $D$ that is important for $M$. Next, we define the explanation views more formally.
We assume an explanation language $\mathbf{EL}$, which is simply a query language that we use as a formalism for defining explanation views. An explanation view $V$ states what component of the database $D$ (existing or derived) is relevant for $M$ to make the prediction $M_{D}(s)$ on any instance $s$; for that, we assume that $V$ is parameterized by $s$, denoted $V\mathord{\langle s\rangle}$. This means that $V\mathord{\langle s\rangle}$ is an ordinary database query with the result $V\mathord{\langle s\rangle}(D)$, for every instance $s$. Recall that $s$ is represented as a key tuple in $D(R)$; hence, to represent explanation views, we assume that queries in $\mathbf{EL}$ can refer to the attributes of $\mathsf{key}(R)$ as constants.
Example 1.
As an example, suppose that our schema consists of $R(A)$ and $T(A,B)$, and $s\in D(R)$. The following explanation view (in SQL) returns the $T$-tuples that connect to $s$:
create view $V\mathord{\langle s\rangle}$ as
select * from $T$
where $T.A=s.A$
Note that the parameterized view $V$ does not necessarily have to refer to its parameter in its definition. If $V$ ignores its parameter $s$, we say that it is instance-agnostic, otherwise it is instance-specific. Formally, $V$ is instance-agnostic if, for each database $D$ and instances $s$ and $s^{\prime}$ it holds that $V\mathord{\langle s\rangle}(D)=V\mathord{\langle s^{\prime}\rangle}(D)$. As an example, the following explanation view is instance-agnostic, as it refers to all tuples of $T$, regardless of their relationship to $s$:
select * from $T$
Finally, by an explanation $E$ we mean a finite subset of concrete queries from $\mathbf{EL}$. The quality of an explanation is measured using two criteria: determinacy (i.e., whether $E$ is enough to determine the behavior of $M$) and conciseness (i.e., how easy it is to grasp $E$).
3.2. Soft Determinacy
Consider an explanation $E$ of an $R$-model $M$ on a database $D$. Intuitively, $E$ “sufficiently determines” $M$ if $M$ behaves similarly even if $D$ changes, as long as the result of $E$ is not affected by this change. Ideally, for any two databases that agree on the explanation views in $E$, the output of model $M$ should be identical. Viewing $M$ as a query, this strict sufficiency notion is precisely determinacy, as described in Section 2.1. To capture the imperfect nature of explanations more effectively, we relax the notion of determinacy in two dimensions. First, we do not require exact determinacy; instead, we account for approximate agreement by quantifying the distance between model predictions before and after the change (see Figure 1). Second, we move away from deterministic determinacy by providing a probabilistic adaptation. For that, we require determinacy in expectation, using our original database $D$ as reference and assuming a probability distribution of contingency databases. Applying the two relaxations leads to an approximate probabilistic notion of determinacy, namely soft determinacy.
Formally, we assume a distribution $\Delta$ over contingency databases that agree on the schema $\mathcal{S}$ of $D$. To be comparable to $M(D)$, we require each database $D^{\prime}$ in this distribution to agree with $D$ on the keys of $R$, that is:
$D^{\prime}(R)[\mathsf{key}(R)]=D(R)[\mathsf{key}(R)]\,.$
In order to define approximate agreement, suppose that $M$ maps an instance $s$ to $a$, and an instance $s^{\prime}$ ito $a^{\prime}$. We use a distance metric dist that gives a quantity $\textrm{dist}(a,a^{\prime})$ of how far $a$ is from $a^{\prime}$. For example, consider the special case where $M$ is a binary classifier with predictions $\in\{0,1\}$; then, it makes sense to use $d$ that checks for identity: $\textrm{dist}(a,a^{\prime})$ is 0 if $a=a^{\prime}$, and 1 otherwise.
Using the above, we arrive at the property of soft determinacy for an explanation $E$, quantified by deviation from determinacy, so that lower deviation implies a better explanation.
Definition 2 (Deviation from Determinacy).
For an instance $s$, and a distribution $\Delta$ of contingency databases, the (instance-specific) deviation from determinacy $\mathsf{dev}_{\Delta}(E,s)$ of an explanation $E$ is the expected distance between the prediction for $s$ on $D$ and the prediction on a contingency $D^{\prime}$, given that $D^{\prime}$ respects the views:
$\mathsf{dev}_{\Delta}(E,s)\vcentcolon=\underset{D^{\prime}\sim\Delta}{\mathbb{E}}\Big{[}$ $\textrm{dist}(M_{D}(s),M_{D^{\prime}}(s))\;\Big{|}\bigwedge_{V\in E}V\mathord{\langle s\rangle}(D)=V\mathord{\langle s\rangle}(D^{\prime})\Big{]}$
Deviation from determinacy of $E$ is then the average over all instances:
$\mathsf{dev}_{\Delta}(E)\vcentcolon=\frac{1}{|D(R)|}\sum_{s\in D(R)[\mathsf{key}(R)]}\hskip-11.38109pt\mathsf{dev}_{\Delta}(E,s).$
3.3. Explanation Conciseness
We associate with each $V\in\mathbf{EL}$ a cost related to its complexity. This cost can be syntactic, e.g., the number of symbols involved in the SQL representation. Alternatively, it can be semantic, e.g., the number of tuples in the view $|V(D)|$. We denote this cost by $\mathsf{cost}(V)$. We then define the cost of an explanation $E$ as $\mathsf{cost}(E)=\sum_{V\in E}\mathsf{cost}(V)$. Note that the syntactic cost does not depend on the database $D$, as the SQL representation refers only to the schema $\mathcal{S}$.
The computational problem of finding the best explanation can be phrased as an optimization problem:
Definition 3 (Soft determinacy explanations).
Given an $R$-model $M$ and a database $D$, the soft determinacy explanation problem asks for a solution $E\subseteq\mathbf{EL}$ to the following optimization objective:
$\min_{E}[\mathsf{dev}_{\Delta}(E)+\lambda\cdot\mathsf{cost}(E)],$
for some trade-off captured by $\lambda\in\mathbb{R},\lambda\geq 0$.
Clearly, there is a natural trade-off between achieving a high level of determinacy (i.e., the minimization of $\mathsf{dev}_{\Delta}(E)$) and low explanation complexity (the minimization of $\mathsf{cost}(E)$). Consider the extreme case of $E$ containing the entire schema $\mathcal{S}$. In this case, $E$ fully determines the model, but also has the complexity of the model that we wish to explain to begin with. On the other hand, a completely uninformative explanation $E=\emptyset$ has minimal complexity, but is likely to hide what the model requires for its predictions.
The computational objective involves the estimation of a conditional probability, a generally intractable task, in order to compute $\mathsf{dev}_{\Delta}(E)$. In Section 4 we discuss several choices for realizing this objective into a tractable optimization problem. Additionally, in Section 5 we propose a learning-based solution for efficient explanation view discovery in the case of GNNs.
4. Instantiating the Framework for Selected SQL Fragments
The framework presented in Section 3 is abstract and requires the choice of an explanation language $\mathbf{EL}$ together with a concrete metric for conciseness and the distribution of contingency databases $\Delta$. In this section, we give specific instantiations that we deem useful and general. In the next sections, we propose heuristic implementations of these over GNN models and test them in experiments.
4.1. Explanation Languages
Our explanation framework highlights the database components that provide the most useful information for the model. These may include important features, important feature combinations, or influential values and value ranges of particular features. In SQL terms, these explanations can be expressed through three core operations: projecting only the relevant attributes of relations, joining relations to reveal informative combinations, and selecting tuples that satisfy specific conditions. To capture these explanations, we focus on three simple SQL fragments utilizing the fundamental relational algebra operations: projections, joins, and selections. Throughout the remainder of this paper, we restrict our attention to instance-agnostic explanation views, for which we define suitable choices of $\mathbf{EL}$ and later propose efficient algorithms. We leave the study on instance-specific explanations for future work.
To define each $\mathbf{EL}$ precisely, we classify the attributes of a relation schema $\mathcal{R}$ in a database schema $\mathcal{S}$ into three categories. Given $\mathsf{att}(\mathcal{R})=(A_{1},\dots,A_{n})$, an attribute $A_{i}$ is a key attribute if it appears in the key signature $\mathsf{key}(\mathcal{R})$. It is a foreign-key attribute if it appears on the left-hand side of a foreign-key constraint in $\mathsf{FK}(\mathcal{S})$. Any attribute that is neither a key nor a foreign key is called a data attribute. In the definitions of $\mathbf{EL}$ that follow, these three attribute types are treated separately.
(1) The language Projection can be used to highlight subsets of data attributes corresponding to important features, while always preserving all key attributes and foreign-key attributes. Formally, Projection contains views of the form:
select $\mathsf{key}(R)$, $\mathord{\mathbf{F}}$, $A_{1},\dots,A_{k}$ from $R$
which project onto a chosen subset of data attributes $A_{1},\dots,A_{k}$, while retaining the sequence $\mathsf{key}(R)$ of key attributes and the sequence $\mathord{\mathbf{F}}$ of foreign-key attributes of $\mathcal{S}(R)$.
(2) The language FKJoin consists of foreign-key joins. Each explanation view in FKJoin is associated with a foreign-key constraint $R[\mathord{\mathbf{C}}]\sqsubseteq\mathsf{key}(S)$ in the schema, and has the form:
select * from $R$ as $r$, $S$ as $s$ where $r[\mathord{\mathbf{C}}]=s[\mathsf{key}(S)]$
These views retain all tuples produced by joining $R$ and $S$ using foreign-key references. Note that more complex join patterns can be captured by multiple views in an explanation view set. Because FKJoin focuses on foreign-key attributes, it always preserves all key attributes as well as the data attributes of relations that appear in at least one explanation view.
(3) The language Selection focuses on filtering tuples based on attribute values. As with Projection, the emphasis is on data attributes, while key and foreign-key attributes are always preserved. Formally, Selection contains views of the form:
select * from $R$ where $\phi$
where $\phi$ is a selection predicate that identifies an informative subset of tuples. We assume that Selection does not affect the graph topology, hence we implicitly assume that an explanation in Selection includes all views of the form “select $\mathsf{key}(R)$, $\mathord{\mathbf{F}}$ from RR” for every $R\in\mathsf{rel}(\mathcal{S})$, where $\mathsf{key}(R)$ contains all key attributes and $\mathord{\mathbf{F}}$ contains all foreign-key attributes of $\mathcal{S}(R)$.
For example, consider a database with a relation Product $P$. The explanation view
select * from $P$ where $P.price<50$ and $P.cat=$ “electronics”
indicates that low-priced electronics constitute an informative region of the data for the model.
Cost model
Explanation conciseness requires a precise definition of the cost model so that every explanation view $V\in E$ is associated with a value $\mathsf{cost}(V)$. Here we propose a simple cost model that focuses on minimizing the description cost of explanation views, making them easier for a user to understand. For Projection, $\mathsf{cost}(V)$ is the number of highlighted data attributes, i.e., $\mathsf{cost}(V)=k$. For FKJoin, each binary join incurs a cost of $1$, i.e., $\mathsf{cost}(V)=1$. For Selection, $\mathsf{cost}(V)$ depends on the complexity of the selection predicate $\phi$. Later, we will consider disjunctions of the form $\phi_{1}$ or $\dots$ or $\phi_{l}$, where each $\phi_{i}$ is an atomic predicate. In this case we simply assign $\mathsf{cost}(V)=l$, i.e., the number of atomic predicates involved.
Language composition
We consider combinations of the three languages by constructing composite explanation views in the relational algebra sense. Importantly, these combination languages produce composite views rather than unions of separate explanation views belonging to one of the three defined languages. For composite views, the cost model is naturally extended: we sum the number of data attributes in projections, the number of joins, and the number of selection predicates involved.
4.2. Database Perturbations
Recall that the definition of deviation from determinacy (Definition 2) allows the liberty to choose any distribution $\Delta$ of contingency databases, as long as they agree with $D$ on the explanation views $E$.
Let us consider two extreme scenarios to illustrate the impact of the choice of $\Delta$. First, consider a $\Delta$ over databases $D^{\prime}$ that differ from $D$ in just one value of a single tuple. A well-behaved model would likely give a similar output on $D$ and $D^{\prime}$, irrespective of the chosen explanation $E$. Therefore, such $\Delta$ would not yield a discriminative criterion between candidate explanations. On the other hand, suppose $\Delta$ is the uniform distribution over all possible $D^{\prime}$, no matter how different they are from $D$. In this case, it is not clear how one could obtain samples from $\Delta$ and, more importantly, the sampled $D^{\prime}$ could be unrepresentative for the application domain of the original database $D$. Ideally, we would like to limit the explanatory power of $E$ to the useful regime, i.e., restricting $\Delta$ to a local neighborhood of $D$ that captures realistic variations while preserving relevance to the application domain.
For each explanation language $\mathbf{EL}$, we give a number of strategies for realizing $\Delta$ by perturbing the original database $D$. The resulting databases $D^{\prime}$ are substantially different from $D$, yet maintain certain properties of $D$, achieving a balance in the aforementioned tradeoff.
(1) Projection
The core idea is to permute the values of columns that are not used in any projection explanation, inspired by a well-known model-agnostic feature selection technique, namely permutation feature importance (PFI) (Breiman, 2001; Fisher et al., 2019; Molnar, 2025). Such column-wise permutations respect the active domain of each attribute and preserve aggregate statistics within a column, while breaking correlations between a permuted column and the model output.
Example 1.
The example in Figure 2 shows two column-wise permutation strategies. The attributes $A$ and $B$ are part of the explanation $E$, whereas $C$ and $F$ are not. Thus, we permute $C$ and $F$ to produce different $D^{\prime}$. We can either apply the permutations to both attributes $C$ and $F$ jointly (left) or independently (right). Note that the latter completely breaks correlations between $C$ and $F$, possibly creating unnatural tuples in $D^{\prime}$. In this example, we have $F=2\cdot C$ for all tuples in $D$. Joint permutation preserves this property in $D^{\prime}$.
Formally, let $\mathsf{Attr}(E)$ be the set of attributes used in views of projection explanations $E$. Recall that $\mathsf{Attr}(E)$ includes all key attributes. Because explanation views must be respected by $\Delta$, we can only permute attributes not present in projection explanations. For each relation name $R\in\mathsf{rel}(\mathcal{S})$, let $\mathsf{attPerm}(R)$ be the set of attributes in $\mathsf{att}(R)$ not in $\mathsf{Attr}(E)$, which are allowed to be permuted.
Let $r=D(R)$. A relation $r^{\prime}$ is a column-wise permutation of $r$ under Projection explanation $E$, written $r^{\prime}\in\mathsf{perm}_{E}(r)$, if $r^{\prime}$ has the same schema and cardinality as $r$, and for each tuple $t^{\prime}\in r^{\prime}$ there exists a unique tuple $t\in r$ with $t^{\prime}[\mathsf{key}(R)]=t[\mathsf{key}(R)]$ such that for all $A\in\mathsf{attPerm}(R)$, $t^{\prime}[A]$ is a permutation of the multiset of values of the attribute $A$ in $r$, while for $A\notin\mathsf{attPerm}(R)$, $t^{\prime}[A]=t[A]$. The joint column-wise permutation $r^{\prime}$ of $r$, written $r^{\prime}\in\mathsf{jointperm}_{E}(r)$, is defined similarly, but all attributes in $\mathsf{attPerm}(R)$ are permuted by the same permutation, preserving $t^{\prime}[\mathsf{key}(R)]=t[\mathsf{key}(R)]$ for the unique corresponding $t\in r$.
The column-wise permutation distribution $\Delta^{\mathsf{ind}}_{E}$ is the uniform distribution with support: $\{D^{\prime}\mid\forall R\in\mathsf{rel}(\mathcal{S}):D^{\prime}(R)\in\mathsf{perm}_{E}(D(R))\}$. The joint column-wise permutation distribution $\Delta^{\mathsf{joint}}_{E}$ is the uniform distribution with support: $\{D^{\prime}\mid\forall R\in\mathsf{rel}(\mathcal{S}):D^{\prime}(R)\in\mathsf{jointperm}_{E}(D(R))\}$. In Section 6 we conduct experiments with both $\Delta^{\mathsf{ind}}_{E}$ and $\Delta^{\mathsf{joint}}_{E}$ and show that the deviation from determinacy is robust to this choice.
Figure 2. Example permutations for Projection (upper) and Selection (lower). For Projection, only attributes $A$ and $B$ (in orange) are in $\mathsf{Attr}(E)$ of explanation $E$. For Selection, the tuples (in orange) $\{(1,0,1,2),(2,1,2,4)\}$ are in $\mathsf{Tups}(E)$. The key attribute $A$ is always retained.
(2) FKJoin
In the case of FKJoin, the goal of a perturbation is to make joins uninformative unless they appear in $E$. Since the joins we consider always correspond foreign-key constraints, we apply a perturbation on the foreign key side. This perturbs the target join without affecting other joins on the side of the primary key.
Example 2.
The example in Figure 3 (left) shows perturbations applied to the primary-foreign pair $T_{2}.B$ - $T_{1}.B$. One could consider permuting the foreign key $T_{1}.B$ as in a projection explanation. But this strategy would not be effective in breaking the predictive signal encoded in the actual join condition as it would retain the number of primary-foreign key pairs per unique value of $T_{1}.B$, i.e., $1$ would appear 4 times while $2$ and $3$ would each appear once. In order to break the join information, we randomly replace each value of $T_{1}.B$ with a value in the domain $\{1,2,3\}$, either uniformly at random (right) or mimicking the underlying frequency distribution of foreign keys (middle). In the former each value in $\{1,2,3\}$ appears $2$ times while in the latter $2$ appears 4 times while $1$ and $3$ each appear once, preserving the frequencies $4,1,1$.
Formally, let $\mathsf{FK}(E)$ denote the set of foreign keys that appear in the explanation $E$. For each $R\in\mathsf{rel}(\mathcal{S})$ and $r=D(R)$, we define a perturbed relation $r^{\prime}$ by independently replacing the values of foreign-key attributes outside of $\mathsf{FK}(E)$ with other values from their respective domains. We consider two alternative strategies for generating $r^{\prime}$. In the uniform replacement $\Delta^{\mathsf{uniform}}_{E}$, each foreign key value is replaced independently and uniformly at random from its domain. In the frequency-preserving replacement $\Delta^{\mathsf{freq}}_{E}$, the multiset of foreign key values in the perturbed relation matches the multiset of frequencies in the original relation: if a foreign key value occurs $i$ times in $r$, then there exists some foreign key value in $r^{\prime}$ that occurs $i$ times, although not necessarily in the same tuples.
Figure 3. Example 2: foreign key perturbations for FKJoin. Here, the foreign key $T_{1}.B$ is perturbed as it is is not in $\mathsf{FK}(E)$.
(3) Selection
For Selection we adapt the column-wise permutation strategies used in the case of Projection. Here, we permute all data attributes only for the tuples that do not satisfy the selection condition of the explanation view. We provide an example in Figure 2 (lower). Specifically, let $\mathsf{Tups}(E)\subseteq\bigcup\{D(R)|R\in\mathsf{rel}(\mathcal{S})\}$ be the set of tuples that satisfy the selection condition. Each relation $r=D(R)$ is separated into $r_{E}$ and $r_{E}^{\prime}$ such that $r_{E}\subseteq\mathsf{Tups}(E)$ and $r_{E}^{\prime}=r\setminus r_{E}$. We then apply column-wise permutations, either independently or jointly, only in $r_{E}^{\prime}$ with $\mathsf{Attr}(E)$ including all data attributes, i.e., only preserving key and foreign key attributes.
5. GNN-Specific Approach
In this section, we focus on models that are based on hetero-GNNs, as defined in Section 2, and propose a heuristic approach to solving the optimization in Definition 3. One approach to face this challenge is to exhaustively search over the space of candidate explanation views and estimate the deviation from determinacy of each candidate. However, such direct optimization is intractable, as the space of possible views is prohibitively large. As an alternative, we leverage the fact that deep learning libraries are naturally equipped with gradient-based continuous optimization; we use this optimization to efficiently discover locally optimal explanation views via a surrogate continuous optimization problem.
5.1. Mask Learning: From Discrete to Continuous Optimization
To formulate a continuous optimization problem, we adopt mask learning, a well-established technique for model interpretability (Fong and Vedaldi, 2017; Ying et al., 2019). Mask learning introduces real-valued learnable masks that can then be mapped to candidate explanations in $\mathbf{EL}$, as we will see shortly. In general, mask learning leverages the fact that deep learning models operate with internal feature representations. Masking means selectively hiding some features while retaining others, ideally keeping only the most important ones. [^1] Notice that we use the term masking in the common ML “selector” convention where higher mask values reveal more of the original signal. Concretely, let $\mathord{\mathbf{x}}\in\mathbb{R}^{d}$ be a feature vector. Applying a mask component $m\in[0,1]$ to $\mathord{\mathbf{x}}$ means replacing $\mathord{\mathbf{x}}$ with a new feature vector as follows:
$\mu(\mathord{\mathbf{x}},m)=m\cdot\mathord{\mathbf{x}}+(1-m)\cdot\mathord{\mathbf{u}},$
where $\mathord{\mathbf{u}}$ is a replacement vector of the same dimension as $\mathord{\mathbf{x}}$. If $m=0$, this replaces $\mathord{\mathbf{x}}$ with $\mathord{\mathbf{u}}$. In addition, if $\mathord{\mathbf{u}}$ is uninformative (e.g., the zero vector), masking removes the influence of $\mathord{\mathbf{x}}$. Multiple mask components can be applied in the model simultaneously on different feature vectors $\mathord{\mathbf{x}}$. The collection of these mask components is a mask $\mathord{\mathbf{m}}$. For any value of the mask $\mathord{\mathbf{m}}$, we can obtain a model output, enabling mask components to be learned via gradient descent with an appropriate loss function.
5.2. Language-Specific Masking
[…]
Figure 4. RDL model pipeline with hetero-GNNs. The figure shows where different masks are located in the pipeline.
Depending on $\mathbf{EL}$, masking is applied at the feature encoding stage or the message-passing stage of the model, as depicted in Figure 4.
(1) Projection
For projection explanations, we apply mask components attribute-wise in the feature encoding space. Since projection explanations necessarily include all primary and foreign keys, we restrict masks to non-key attributes.
For each relation $R\in\mathcal{S}$, we assign one mask component $m_{R.A_{i}}\in[0,1]$ per attribute $A_{i}\in\mathsf{att}(R)\setminus\big{(}\mathsf{key}(R)\cup\mathsf{FK}(R)\big{)}$. For each node $v=(R,t)$, the same mask vector $(m_{R.A_{1}},\ldots,m_{R.A_{k}})$ is applied at feature encoding, changing Equation 2 as follows:
$\begin{split}\mathord{\mathbf{X}}_{v}=\mathsf{Enc}_{R}(&\mu(\mathsf{Enc}_{R.A_{1}}(t[A_{1}]),m_{R.A_{1}})\oplus\cdots\oplus\\ &\mu(\mathrm{Enc}_{R.A_{k}}(t[A_{k}]),m_{R.A_{k}}))\end{split}$
Recall that $\mu(\mathord{\mathbf{x}},m)$ applies the mask component $m$ to the encoding vector $\mathord{\mathbf{x}}$. When masking attribute $A_{i}$, we use a replacement vector $\mathord{\mathbf{u}}=\mathsf{Enc}_{R.A_{i}}(t^{\prime}[A_{i}])$ for a uniformly and randomly sampled tuple $t^{\prime}\in D(R)$. This ensures the replacement of the attribute’s encoding is uninformative, yet realistic. Masks for all relations and attributes are learned jointly.
After mask learning, continuous mask values are mapped into projection explanations. For that, we use a threshold $\delta>0$:
$m^{\prime}_{R.A_{i}}=\begin{cases}1&\text{if }m_{R.A_{i}}\geq\delta\\ 0&\text{otherwise}\end{cases}$
Attributes with $m^{\prime}_{R.A_{i}}=1$ are included in the projection explanations; the rest are omitted.
(2) FKJoin
For join explanations, we apply masks during message-passing. Every primary/foreign-key pair corresponds to an edge type in the graph (see Section 2.4). Thus, we learn a mask value $m_{\lambda_{i}}\in[0,1]$ for every edge type $\lambda_{i}$ in order to decide whether to retain the corresponding join in the explanation. Equation 1 is adapted as follows:
$\begin{split}\mathord{\mathbf{h}}^{(\ell)}(v)=\mathsf{UPD}(\mathord{\mathbf{h}}^{(\ell-1)}(v)\,,\,&\mu(\mathcal{A}^{(\ell-1)}(v,\lambda_{1},o_{1}),m_{\lambda_{1}}),\dots,\\ &\mu(\mathcal{A}^{(\ell-1)}(v,\lambda_{q},o_{q}),m_{\lambda_{q}}))\end{split}$
Notice that for each edge type $\lambda_{i}$, the same mask component $m_{\lambda_{i}}$ is applied at the aggregate vector of messages irrespective of the direction $o_{i}$, as both directions correspond to the same join. Also, the same mask components are used across all layers $\ell$, because message-passing corresponds to the same join operation irrespective of the GNN layer. We replace each aggregate with a zero replacement vector $\mathord{\mathbf{u}}$. Similarly to Projection, continuous mask values are mapped to join explanations after thresholding.
(3) Selection
First, consider the case of a single selection predicate $\phi$. We assign a mask $m_{\phi}\in[0,1]$ to all tuples that satisfy $\phi$. For each node $v=(R,t)$ with $\phi(t)=1$, we set $m_{(R,t)}=m_{\phi}$. Masking is then applied per tuple by adapting Equation 2 into:
$\mathord{\mathbf{X}}_{v}=\mu(\mathsf{Enc}_{R}\left(\mathsf{Enc}_{R.A_{1}}(t[A_{1}])\oplus\cdots\oplus\mathrm{Enc}_{R.A_{k}}(t[A_{k}])\right),m_{(R,t)})$
Similarly to Projection, when masking $v=(R,t)$, we replace $\mathord{\mathbf{X}}_{v}$ with $\mathord{\mathbf{u}}=\mathord{\mathbf{X}}_{v^{\prime}}$ for a randomly sampled $v^{\prime}=(R,t^{\prime})$, where $t^{\prime}\in D(R)$.
For the implementation of Selection, we consider conditions that involve disjunctions of atomic predicates $\phi_{1}\vee\dots\vee\phi_{l}$. Following real-valued logic techniques that map logical formulas into differentiable form (Li et al., 2019), we employ the Łukasiewicz t-conorm for disjunctions. Thus, the mask values per tuple are calculated as $m_{(R,t)}=\min\!\left\{1,\sum_{\phi_{i}(t,\mathord{\mathbf{c}}_{i})=1}m_{\phi_{i}}\right\}$. A similar strategy may be applied for conjunctions. After learning masks for all candidate selection predicates jointly, we apply thresholding to map them into discrete explanations.
To combine selections and projections in a composite language, namely ProjSelect, we first learn projection masks as discussed for Projection, and then learn masks for candidate selection predicates that use data attributes present in the learned projections. In our experiments (Section 6), we restrict ourselves to categorical attributes with equality predicates and numerical attributes with range predicates defined via quantiles.
5.3. Mask Optimization
We define the loss function for mask learning, a surrogate optimization for our original problem in Definition 3:
$\min_{\mathord{\mathbf{m}}}\ \mathcal{L}_{\text{task}}+\lambda\,\mathcal{L}_{\text{reg}}(\mathord{\mathbf{m}}),$
where the mask $\mathord{\mathbf{m}}$ is a vector that contains all mask components to be learned, $\mathcal{L}_{\text{task}}$ is the original loss of the model (e.g., classification or regression loss) and $\mathcal{L}_{\text{reg}}$ is an $\ell_{1}$-regularization term applied to $\mathord{\mathbf{m}}$ that minimizes the size of the mask. Note that $\mathcal{L}_{\text{task}}$ ensures that the learned masks preserve the model’s behavior, a goal that is aligned with the soft determinacy objective of Definition 3. On the other hand, the $\ell_{1}$ penalty encourages sparsity in the masks (Hastie et al., [n.d.]), promoting concise explanations by selecting only the most relevant database components.
6. Experimental Evaluation
We assess the performance of our proposed method and the overall effectiveness of the framework.
Metrics
The main metric we consider is deviation from determinacy $\mathsf{dev}_{\Delta}(E)$ (see Definition 2). The distance dist is the absolute difference for binary classification and the normalized absolute difference $\textrm{dist}(x,y)=\frac{|x-y|}{|x|+|y|}$ for regression, both bounded in $[0,1]$. Each $\mathsf{dev}_{\Delta}(E)$ estimate is averaged over 5 perturbation samples; we report both mean and standard deviation. Second, we evaluate the explanation size $k$. Recall that this is the number of projected data attributes for Projection, the number of join conditions for FKJoin, and the number of predicates for Selection. Finally, we report the time to generate the explanation, which corresponds to training time for mask-based methods.
Data
We use RelBench, the standard benchmark suite for RDL (Robinson et al., 2024), which includes real-world relational databases and diverse predictive tasks. We focus on node-level tasks, i.e., regression and binary classification. Details of the datasets are provided in the online appendix. As Table 2 shows, we denote each combination of database $i$ and task $j$ as DiiTjj.
Models to explain
We train RDL models from scratch. For each database, we construct a graph representation (as explained in Section 2.4) and use PyTorch Frame (Hu et al., 2024) to implement feature encoders. The GNN is implemented in PyG (Fey et al., 2025), with 32-dimensional encoders. We perform hyperparameter tuning over the learning rate, batch size, and number of GNN layers. Models are trained using the train/validation/test split provided by the benchmark. Our models reach performance similar to the one previously reported (Robinson et al., 2024) for GNN approaches. For each explanation task, we sample 100 instances for mask training and 100 instances for explanation evaluation. Since many classification tasks are imbalanced, we use balanced sampling to ensure that the selected instances contain an equal number from each class. For temporal tasks we perform temporal sampling (Fey et al., 2024) to ensure time-consistency.
Methods
1
Mask-based: We implement the learnable masks as introduced in Section 5, with early stopping. 1a Column Mask: masks for Projection. 1b FKPK Mask: masks for FKJoin. 1c Filter Mask: masks for Selection.
Since no prior work generates SQL explanations for RDL, we implement a set of baselines. 2 Ranking-based for Projection: We implement two baselines that order data attributes according to their importance, and select the top-$k$ for each target explanation size $k$. A known limitation of such methods is that they ignore dependencies between attributes (Molnar, 2025). 2a Local Impact: The importance of a data attribute is computed as $dev_{\Delta}$ when the attribute is taken as an explanation by itself. The lower the $dev_{\Delta}$ the more important the attribute. 2b PFI: We adapt Permutation Feature Importance (PFI) (Molnar, 2025) by measuring the change in $dev_{\Delta}$ when each attribute is removed. Attributes are ranked by impact, where higher deviation implies higher importance.
3
Greedy approach for Projection and FKJoin: 3a Greedy: Starting from the empty set, we iteratively add the attribute that yields the largest reduction in $dev_{\Delta}$. 3b Greedy Expansion: At each step, we add a join that maximally reduces $dev_{\Delta}$ while ensuring connectivity to the prediction entity. This is more computationally feasible than Greedydue to fewer join candidates at each iteration.
5
Random Subset for Projection: We randomly sample the data attributes and report the average over 5 samples.
4
Empty: An explanation containing no information, resulting in a full database perturbation. We use this as a point of reference.
Implementation details
For the embeddings of textual attributes we use a BERT-based model [^2] https://huggingface.co/sentence-transformers/distilbert-base-nli-mean-tokens. All experiments are conducted on a VM with an RTX 6000 Ada GPU, 14 vCPUs, and 188 GB of RAM. Our code is available at https://github.com/agapiR/rdl-explain.
Evaluation plan
We evaluate the proposed mask learning strategy separately for the Projection and FKJoin languages. Our goal is to evaluate the tradeoff between conciseness and soft determinacy. To ensure a fair comparison, we use the same explanation size for all methods, denoted as $k^{*}$, which we determine by thresholding the learned mask as discussed in Section 5. We also assess the robustness of $dev_{\Delta}$ under different database perturbation strategies $\Delta$. Additionally, we conduct a case study on one task where we showcase Selection and $3$ composite languages.
6.1. Projection Evaluation
Figure 5 reports the results for Projection across all datasets and tasks, including averages (AVG).Figure 5(a) and Figure 5(b) show $\mathsf{dev}_{\Delta}$ for two different perturbation strategies. The main takeaway is that Column Mask produces higher quality explanations (in terms of $\mathsf{dev}_{\Delta}$) than the baselines on most tasks, and on average, while outperforming them on execution time (Figure 5(c)) by 1-2 orders of magnitude for large databases. This is expected since Column Mask’s running time is dependent on the size of the model and the size of the explanation task, which is usually small. Among the baselines, PFI gives the best results both on quality and on time. However, all baselines suffer from their costly dependence on the schema size, i.e., the number of data attributes that they consider. On database D11 PFI is faster, but this is due to the fact that the database is small and has a concise schema. Comparing the two perturbation strategies, they appear to yield very similar results, indicating that $\mathsf{dev}_{\Delta}$ is robust to the choice of permutation strategy.
While Figure 5 assumes the same explanation size $k^{*}$, Figure 6 shows $dev_{\Delta}$ for various explanation sizes $k$, illustrating the tradeoff between the two. As the size $k$ increases, we generally observe $dev_{\Delta}$ decreasing, i.e., the explanation quality increasing. This is not true for Random Subset, which shows that the choice of attributes in the explanation is crucial. For smaller $k$ than $k^{*}$, the baselines are often better than Column Mask, but Column Mask usually outperforms them for the optimal $k^{*}$ (i.e., the end of the axis).
(a) Deviation from determinacy $\mathsf{dev}_{\Delta}$ for each database-task and database perturbation $\Delta^{\mathsf{ind}}_{E}$.
(b) Deviation from determinacy $\mathsf{dev}_{\Delta}$ for each database-task and database perturbation $\Delta^{\mathsf{joint}}_{E}$.
(c) Time to obtain explanation for each database-task and $\Delta^{\mathsf{ind}}_{E}$. Results for $\Delta^{\mathsf{joint}}_{E}$ are similar.
Figure 5. Evaluation for Projection. For visualization purposes, tasks are separated into “easy” (left) and “hard” (right) ones, with separate $\mathsf{dev}_{\Delta}$ scales. The average (AVG) across all tasks is also shown. Below the dataset and task name, we indicate the explanation size $k^{*}$ and the total number of data attributes.
(a) D1T1
(b) D1T2
(c) D1T3
(d) D6T1
(e) D6T2
(f) D6T3
(g) D4T1
(h) D4T2
Figure 6. Soft determinacy versus conciseness for Projection explanations, for 66 tasks and database perturbation $\Delta^{\mathsf{ind}}_{E}$.
6.2. FKJoin Evaluation
Figure 7 shows the corresponding set of results for FKJoin. As in Projection, the mask-based approach FKPK Mask is significantly faster than the greedy baseline Greedy Expansion. In terms of deviation from determinacy, the gap between FKPK Mask and Greedy Expansion is smaller, but FKPK Mask still edges out on average. The differences between the two perturbations strategies are again minor, verifying the robustness of our metric.
(a) Deviation from determinacy $\mathsf{dev}_{\Delta}$ for each database-task and database perturbation $\Delta^{\mathsf{uniform}}_{E}$.
(b) Deviation from determinacy $\mathsf{dev}_{\Delta}$ for each database-task and database perturbation $\Delta^{\mathsf{freq}}_{E}$.
(c) Time to obtain explanation for each database-task and $\Delta^{\mathsf{uniform}}_{E}$. Results for $\Delta^{\mathsf{freq}}_{E}$ are similar.
Figure 7. Evaluation for FKJoin. For visualization purposes, tasks are separated into “easy” (left) and “hard” (right) ones, with separate $\mathsf{dev}_{\Delta}$ scales. The average (AVG) across all tasks is also shown. Below the dataset and task name, we indicate the explanation size $k^{*}$ and the total number of foreign-key constraints.
6.3. Case Study
We perform a case study on the rel-trial dataset (Initiative et al., 2019), a large-scale relational dataset from clinical trial reports. It contains 15 tables with a total of 5,852,157 rows and 140 attributes, out of which 110 are data attributes and 30 are key or foreign-key attributes. We focus on the study-outcome task (D6T1), a binary classification task that predicts whether a trial achieves its primary outcome.
Table 1. Case study (Section 6.3): Explanation Metrics by Explanation Language for D6T1.
$\mathbf{EL}$ $\mathsf{dev}_{\Delta}(E)$ SQL Size Data (%)
Projection 0.08 15 11.07
FKJoin 0.03 10 58.43
Selection 0.05 5 86.36
ProjSelect 0.09 20 8.08
FKJoinProj 0.08 25 11.07
FKJoinProjSelect 0.09 30 8.08
Empty 0.25 0 0.0
Table 1 illustrates how simple explanation views can yield low $\mathsf{dev}_{\Delta}$, below 0.1 for compact languages such as FKJoin and Projection. The results also highlight the flexibility offered by the $\mathbf{EL}$ choice. By combining projections, joins, and selections, one can balance size, data reduction, and expressivity.
SELECT id, nct_id,
allocation, intervention_model, primary_purpose
FROM designs d
WHERE d.allocation=0 OR d.intervention_model=0
(a) Example explanation view using ProjSelect.
SELECT * FROM studies s
JOIN facilities_studies fs ON s.nct_id=fs.nct_id
JOIN facilities f ON fs.facility_id=f.facility_id
(b) Example explanation view using FKJoin.
We also qualitatively examine the selection predicates learned by our Filter Mask approach. Our method detects the following predictive selection predicates: designs.allocation = 0 selects single-arm trials, designs.intervention_model = 0 selects trials applying an intervention to a single group and the predicate studies.phase $\in\{0,2,4\}$ selects trials based on the phase. Figures 8(a) and 8(b) show two example explanation views. The table designs originally contains $15$ attributes, of which we maintain only $3$ data attributes, together with the key and foreign-key attributes. In Figure 8(a), we apply the disjunction of two selection predicates which together select the study designs that correspond to single-arm clinical trials or single-group intervention trials. We can infer that these trials might be simpler, cheaper or faster to conduct. The view indicates that this knowledge about the study’s design is highly predictive of its outcome. In Figure 8(b), the view captures an important join path of length 2. (The two joins correspond to two separate views in our language FKJoin, but they are shown in a single view here.) It indicates that the (potentially multiple) facilities where a study is conducted are important for its outcome.
6.4. Retraining Results
As a sanity check, we retrain the models using only the subset of the data retained by our Projection and FKJoin explanations. As Table 2 shows, performance is comparable for most tasks while the database size is substantially smaller. This confirms that the selected data retain the core predictive signal.
7. Related Work
Learning over a relational database
Statistical relational learning (Getoor and Taskar, 2007; Raedt et al., 2016) is one of the first approaches to directly leverage the relational structure, factorized learning (Schleich et al., 2016; Kumar et al., 2016; Olteanu, 2020) avoids costly joins, automatic feature engineering (Kanter and Veeramachaneni, 2015; Chepurko et al., 2020; Zhang et al., 2023; Stanoev et al., 2024) synthesizes new features, while other approaches train arbitrary (often tree-based (Chen and Guestrin, 2016)) predictive models. The currently emerging approach is Relational Deep Learning (RDL) (Cvitkovic, 2019; Fey et al., 2024; Wang et al., 2024). It models the relational database as a heterogeneous graph, and employs tabular feature encoders (Hu et al., 2024) trained jointly with Graph Neural Networks (GNNs) to solve a wide range of predictive tasks. The current research frontier further builds specialized architectures for RDL that fall under the same GNN paradigm (Hilprecht et al., 2023; Chen et al., 2025; Yuan et al., 2025) and adapts Graph Transformers to the relational setting (Rampášek et al., 2022; Kong et al., 2023; Dwivedi et al., 2025a). Of course, the RDL approach inherits the opacity of “black-box” deep learning, creating the need for explainability. Since RDL models, irrespective of their implementation, make predictions at the database level, we consider explanations that refer to the database rather than the specific modeling components, such as the heterogeneous graph.
ML explanations
To address the explainability in RDL, we review the broader literature on ML explanations. These are categorized by their perspective as either local or global (Guidotti et al., 2018; Samek and Müller, 2019; Molnar, 2025; Bassan et al., 2024). Global approaches aim to explain a model’s overall behavior, while local approaches focus on individual predictions. Approaches also vary in focus: some rely on producing local approximations using simpler models (e.g., LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017)), while others highlight critical input subsets, such as the rule-based conditions generated by Anchors (Ribeiro et al., 2018). A more formal distinction further classifies the latter into abductive and contrastive (Marques-Silva and Ignatiev, 2022; Barceló et al., 2025). Abductive explanations identify minimal subsets of the input sufficient to preserve the model’s prediction, whereas contrastive explanations specify the minimal changes required to alter it. In our context, explanation views are global, as they refer to the database as a whole rather than particular instances, and abductive, as they identify the database components sufficient to determine the model’s behavior.
Subgraph-based GNN explanations
Instance-level explanations specific to Graph Neural Networks (GNNs) aim to answer questions of the type “Why does the GNN produce a particular output for a given instance (i.e., node in the graph)?”. Most of these explanations are at the level of individual graph elements, i.e., the explanation is a subgraph of the input graph. Some instance-level explainers focus on neighboring nodes and their individual features (Pope et al., 2019; Vu and Thai, 2020; Funke et al., 2021; Li et al., 2023a), others on important edges (Ying et al., 2019; Lucic et al., 2022), while others on connected subgraphs (Yuan et al., 2021; Lin et al., 2021; Shan et al., 2021; Li et al., 2023b; Mika et al., 2023). Few are tailored to hetero-GNNs (Li et al., 2023a; Mika et al., 2023) by specializing subgraph explanations to meta-paths. Similarly, global GNN explainers, which provide insights for a GNN across all instances, use subgraphs as explanations. More specifically, some recover a collection of subgraphs that collectively explain many instances (Yuan et al., 2020; Lv and Chen, 2023; Wang and Shen, 2023; Huang et al., 2023; Chen et al., 2024) aiming at explanation generality, while others extract instance-level explanation subgraphs which are then combined into global insights (Lv et al., 2022; Azzolin et al., 2023; Chen et al., 2024). In contrast to our approach, most global explainers implicitly assume the predictive tasks are motif-based (Azzolin et al., 2025b) and they are often evaluated against a ground truth. We refer the reader to a recent survey for more details on these techniques (Yuan et al., 2022). We depart from all these subgraph-based approaches in the following ways: 1 Our explanations are expressed in terms of the relational database, which is the input object, instead of the graph that is internally constructed by the model. 2 The complexity of our explanations is measured in terms of the complexity of SQL views, which can succinctly refer to billions of data points. This implies that subgraph-based explanations can be expressed as SQL views, often much more succinctly than in graph-based representations. 3 Our SQL explanations can express database components that go beyond simple subgraphs. For example, a single-column projection does not correspond to a connected subgraph. It is also possible to provide explanations that correspond to derived information, such as aggregates which do not explicitly exist in the graph at all.
Other GNN explanations
Beyond subgraphs, there are other types of explanations that are fundamentally different. First, some approaches, if adapted to RDL, would not correspond to any database elements (Vu and Thai, 2020; Zhang et al., 2021; Huang et al., 2022; Baldassarre and Azizpour, 2019; Schnake et al., 2021; Schwarzenberg et al., 2019; Schlichtkrull et al., 2021), e.g., because they refer to GNN layers or construct surrogate models. Second, distillation (Pluska et al., 2024) can map the GNN into a logical classifier. This is similar in spirit to our explanations; however, their base predicates refer to inner features of the model and can quickly become too complex. Finally, self-explainable GNNs (Dai and Wang, 2021) claim to provide simpler GNN models, thus with better interpretability properties. Recently this approach has been adapted to RDL, but it is limited to meta-path selection (Ferrini et al., 2025). Our explanations are given post-hoc on an already trained model, instead of training a simpler model from scratch.
Table 2. Databases, tasks, model performance, and performance after training with only the data in our explanations.
[…]
Binary Classification (ROC-AUC)
rel-f1 driver-dnf D1T1 77.77 73.26 -4.51 51.19%
rel-f1 driver-top3 D1T2 76.77 65.29 -11.48 63.53%
rel-avito user-clicks D2T1 66.22 67.08 0.86 1.93%
rel-avito user-visits D2T2 64.48 63.59 -0.89 8.34%
rel-stack user-engage D3T1 89.92 89.58 -0.34 64.56%
rel-stack user-badge D3T2 88.33 88.11 -0.23 84.05%
rel-hm user-churn D4T1 68.78 68.80 0.02 50.84%
rel-trial study-outcome D6T1 71.28 69.78 -1.49 79.46%
rel-event user-repeat D5T1 78.76 78.44 -0.32 94.81%
rel-event user-ignore D5T2 79.35 75.22 -4.13 73.21%
Regression (MAE)
rel-f1 driver-position D1T3 3.3552 3.3996 0.0444 51.22%
rel-avito ad-ctr D2T3 0.0401 0.0457 0.0056 11.86%
rel-hm item-sales D4T2 0.0561 0.0567 0.0006 31.35%
rel-trial study-adverse D6T2 46.2990 45.4508 -0.8482 85.23%
rel-trial site-success D6T3 0.3247 0.3746 0.0500 76.02%
rel-event user-attendance D5T3 0.2404 0.2381 -0.0023 94.51%
Success metrics for explanations
Most prior work focuses on sufficiency of explanations, i.e., finding a small subgraph that (by itself) retains the GNN prediction, quantified by a metric called fidelity [^3] Fidelity is sometimes referred to as fidelity minus (Amara et al., 2022) or sufficiency (Longa et al., 2025) (Yuan et al., 2022). Variants of this metric either focus on retaining prediction confidence (Li et al., 2023a) or robustness under random, (Zheng et al., 2024), infinitesimally small (Agarwal et al., 2022), and even adversarial (Fang et al., 2023) perturbations. In addition to sufficiency, some works assess explanation necessity (Amara et al., 2022; Longa et al., 2025), i.e., whether the explanation subgraph is essential to maintain the GNN prediction. Closest to soft determinacy is a fidelity variant called degree of sufficiency (Azzolin et al., 2025a, b) in that it has a probabilistic nature. Compared to all this prior work, our evaluation has the following key differences: 1  Since the framework is general, soft determinacy is parameterized by the explanation language $\mathbf{EL}$. In contrast to subgraph modifications, the perturbations needed for soft determinacy calculation are defined in the database space and must respect $\mathbf{EL}$ to be considered valid. 2  We minimize explanation complexity with respect to query instead of data size. 3  We also focus on the time to obtain an explanation, which has received limited attention in prior work, with few exceptions (Yuan et al., 2021; Lin et al., 2021; Chen et al., 2024).
Explanations in databases
In the context of databases, explainability research targets various phenomena including why specific query results appear or are absent (why-not explanations) (Glavic et al., 2021; Chapman and Jagadish, 2009), outliers (Miao et al., 2019; Wu and Madden, 2013), trends in query outputs (Roy and Suciu, 2014), and anomalies related to performance (Kalmegh et al., 2018; Zhang et al., 2017). These efforts parallel ML explainability in aiming to explain algorithmic outputs via input data and/or computational logic. Explanations in databases are often expressed in terms of individual tuples (Meliou et al., 2010a, b; Salimi et al., 2016; Livshits et al., 2021) or query provenance (Buneman et al., 2001; Suciu et al., 2011; Green and Tannen, 2017). Differently, in our framework, explanations are themselves simple queries, explaining complex queries (deep models).
8. Conclusions and Future Work
We presented a framework for abductive global explanations of machine-learning models over relational databases. Grounded in a probabilistic adaptation of determinacy, our approach enables interpretable and flexible explanations as relational queries of varying fragments, without depending on model internals. We studied the instantiation of the framework on an architecture based on heterogeneous GNNs, and introduced an efficient masking-based heuristic that circumvents the need to compute soft determinacy directly. Our extensive evaluation on the RelBench benchmark demonstrated that our framework achieves high predictive accuracy with low runtime cost across diverse tasks.
Several directions for future work emerge from our framework. A first class of extensions concerns broadening the applicability and expressive power. One natural direction is to develop efficient instantiations for additional model classes, such as attention-based models that gain traction in relational learning (Wang et al., 2025; Hollmann et al., 2025; Dwivedi et al., 2025b). It is also important to develop solutions for other tasks, such as link prediction. Another direction is to enrich the explanation languages by supporting instance-specific explanations and incorporating constructs such as group-aggregate queries and more complex selection conditions. This raises important design questions, such as how to systematically select good candidate predicates.
Beyond these, a fundamental question is how to identify explanation languages that are naturally aligned with a given use case—can we predict which explanation constructs will yield the most useful results for a particular combination of model architecture and database schema? For instance, it may turn out that group-aggregate explanations are especially well-suited to GNNs, as both rely on aggregation of related tuples (e.g., grouping by the same foreign keys). Formalizing and testing such alignments could enable systems that not only allow flexible adaptation to explanation needs, but also guide users toward proper explanation languages.
Appendix A Experimental Details
Table 3 provides a summary for the Relbench Benchmark including dataset statistics and our models’ performance in comparison to SOTA performance.
Table 3. Datasets, tasks, and performance comparison with SOTA (Robinson et al., 2024; Lachi et al., 2025). Database-level statistics are shared across tasks.
Database DiiTjj Task #Tuples #Attr. #Data Attr. Size (MB) Perf. (Ours) Perf. (SOTA)
Binary Classification (ROC-AUC)
rel-f1 D1T1 driver-dnf 97,606 67 45 1.18 0.7777 0.7355
D1T2 driver-top3 0.7677 0.8473
rel-avito D2T1 user-clicks 24,653,915 42 26 443.83 0.6622 0.6590
D2T2 user-visits 0.6448 0.6620
rel-stack D3T1 user-engagement 5,399,818 52 33 1004.46 0.8992 0.9059
D3T2 user-badge 0.8833 0.8886
rel-hm D4T1 user-churn 16,931,173 37 33 179.99 0.6878 0.6998
rel-event D5T1 user-repeat 59,141,053 246 122 440.74 0.7876 0.7777
D5T2 user-ignore 0.7935 0.8398
rel-trial D6T1 study-outcome 5,852,157 140 110 649.79 0.7128 0.7009
Regression (MAE)
rel-f1 D1T3 driver-position 97,606 67 45 1.18 3.3552 4.022
rel-avito D2T3 ad-ctr 24,653,915 42 26 443.83 0.0401 0.0410
rel-hm D4T2 item-sales 16,931,173 37 33 179.99 0.0561 0.0560
rel-event D5T3 user-attendance 59,141,053 246 122 440.74 0.2404 0.2580
rel-trial D6T2 study-adverse 5,852,157 140 110 649.79 46.2990 44.0110
D6T3 site-success 0.3247 0.4000